# Unveiling the Flaws: A Critical Analysis of Initialization Effect on Time Series Anomaly Detection

### v1

@@ 5 Conclusion
This paper critically analyzed the impact of initialization on Time Series Anomaly Detection (TSAD) models. Our experiments revealed that these models are highly sensitive to parameters like window size, seed number, and normalization, leading to significant performance variability. This sensitivity can be exploited to artificially inflate performance results, raising concerns about the robustness and reliability of current TSAD methods. Our findings highlight the necessity for rigorous evaluation protocols and transparent reporting of preprocessing steps to ensure the reliability and fairness of anomaly detection methods.
To advance the field, we recommend that future research focuses on developing more robust models that are less sensitive to initialization parameters. Additionally, standardizing evaluation protocols and ensuring transparency in reporting preprocessing steps are essential to prevent the exploitation of these sensitivities. By addressing these issues, we can develop more reliable and fair anomaly detection methods that are truly beneficial in real-world applications.

### v2

@@ IV Conclusion
This paper critically analyzed the impact of initialization on TSAD models. Our experiments revealed that they are highly sensitive to parameters like window size, seed number, and normalization, leading to significant performance variability. This sensitivity can be exploited to artificially inflate performance results, raising concerns about the robustness and reliability of current TSAD methods. Our findings highlight the necessity for rigorous evaluation protocols and transparent reporting of preprocessing steps to ensure the reliability and fairness of anomaly detection methods. To advance the field, we recommend that future research focuses on developing more robust models that are less sensitive to initialization parameters. Additionally, standardizing evaluation protocols and ensuring transparency in reporting preprocessing steps are essential to prevent the exploitation of these sensitivities.